OpenStudy (monkeyagf):

The cost of 4 cartons of fresh cream is $5.76. What is the cost of 13 such cartons? A. $16.25 B. $18.72 C. $18.76 D. $36.04

OpenStudy (igreen):

Set up a proportion. \(\sf \dfrac{4}{5.76} = \dfrac{13}{x}\) Cross multiply and solve for 'x'.

OpenStudy (igreen):

\(\sf 4 \cdot x =~?\) \(\sf 5.76 \cdot 13 =~?\)

OpenStudy (andrewherpeche):

its B

OpenStudy (monkeyagf):

ok, thx

OpenStudy (andrewherpeche):

5.76/4=1.44*13=18.72

OpenStudy (igreen):

@andrewherpeche Don't give out answers.

OpenStudy (andrewherpeche):

i told her how i did it tho

OpenStudy (monkeyagf):

yea

OpenStudy (andrewherpeche):

thank you

OpenStudy (monkeyagf):

yw

OpenStudy (monkeyagf):

and thank u for helping

OpenStudy (igreen):

You're supposed to guide her so she comes up with the answer herself, not give it out.